# KERMIT: Knowledge Graph Completion of Enhanced Relation Modeling with Inverse Transformation

### v2

@@ Appendix B Additional experiment results
B.1. Additional ablation results
Figure 4. Model performance of text-based methods with the proposed approaches on FB15k-237 dataset. "w/ inv. rel.": with only inverse relations; "w/ pred. desc": with predictive descriptions. The predictive descriptions are collected using phi3-medium.
B.2. Inverse relations
Table 8. Relations and their corresponding inverses within the WN18RR dataset.
@@ B.2. Inverse relations
Table 9. A subset of the raw, normalized and inverse relations within the FB15k-237 dataset.
@@
B.3. Predictive descriptions
Table 10. Triple examples of the WN18RR dataset along with generated predictive descriptions via phi3-medium.
@@ B.3. Predictive descriptions
Table 11. Examples of the FB15k-237 dataset along with generated predictive descriptions via phi3-medium.
@@ Appendix C Additional analysis results
Table 12. Triple examples of the WN18RR dataset along with generated predictive descriptions via three different LLMs.

### v3

@@ Appendix B Additional experiment results
B.1 Additional ablation results
Figure 5: Model performance of text-based methods with the proposed approaches on FB15k-237 dataset. "w/ inv. rel.": with only inverse relations; "w/ pred. desc": with predictive descriptions. The predictive descriptions are collected using phi3-medium.
B.2 Inverse relations
Table 8 presents the relations and their inverses from the WN18RR dataset. For the FB15k-237 dataset, we show results in Table 9.
Table 8: Relations and their corresponding inverses within the WN18RR dataset.
@@ B.2 Inverse relations
Table 9: A subset of the raw, normalized and inverse relations within the FB15k-237 dataset.
@@
B.3 Predictive descriptions
Table 10 and 11 present additional examples of predictive descriptions for the WN18RR and FB15k-237 datasets, respectively.
Table 10: Triple examples of the WN18RR dataset along with generated predictive descriptions via phi3-medium.
@@ B.3 Predictive descriptions
Table 11: Examples of the FB15k-237 dataset along with generated predictive descriptions via phi3-medium.
@@ Appendix C Additional analysis results
The results of further analysis on the selection of different LLMs are shown in Table 12.
Table 12: Triple examples of the WN18RR dataset along with generated predictive descriptions via three different LLMs.